OpenStudy (anonymous):

You throw a baseball straight up into the air with a speed of 24.5 m/s. how long does it take the baseball to reach its highest point? A. 0.4s B. 101s C. 3.2s D. 2.5s

OpenStudy (theeric):

Hi! You want to solve this by using a conceptual trick, and using math to finish it off. Or there might be an easier formula that I don't know. Here's the concept: The ball's vertical velocity is 0 at its highest point. Now, what do you know?

OpenStudy (theeric):

You know the initial and final velocity as you throw it and it gets to its highest point. You also know that it will be pulled down by gravity. You can't calculate the force of gravity, but it doesn't matter - you know the effect of gravity to be gravitational acceleration. So you have that. And you need the time. Good luck! This sounds a little like how you define acceleration... Define acceleration, put it into algebra, and see what you have! A little algebraic manipulation, and you're good to solve it. Take care!

OpenStudy (anonymous):

Vf= Vi + gt solve for t Vi= 24.5 m/s Vf at the peak is always zero g= 9.8 m/s² 0= 24.5 m/s + (9.8m/s²)t -24.5m/s= 9.8 m/s t t= 2.5 Thanks you really helped :)

OpenStudy (theeric):

You're very welcome! :) I agree!

OpenStudy (theeric):

By the way, The equation I thought of is just a rearranged version of the one you found! However, I recommend that you do whatever you're most comfortable with for now! Acceleration is change in velocity over a given time interval: \(\large a=\dfrac{v_f-v_i}t\)